# The Price of Competitive Information Disclosure

### v1

@@ 1 Introduction
Every day, people make critical decisions based on limited information, but often someone else controls what information they see. For example, when job seekers apply to a company, they highlight their best qualities (such as a prestigious degree or a successful project) and downplay less flattering information. The employer then must decide who to hire based on the résumés and references provided. Similarly, in college admissions, students choose which recommendation letters to send or whether to submit optional test scores, and hence curate the pictures that admissions officers see. In both examples, the party presenting the information has an incentive to reveal the information selectively to improve their chances. Such strategic revelation of information is not lying; it is about being selective in revealing the truth.
From the perspective of someone who aims at optimizing societal outcomes, agents revealing information should only be helpful, since the societal optimizer always has the option of ignoring the information revealed. However, quantitatively, to what extent can information revelation improve societal outcomes? We first look at a simple example.
@@
Consider a scenario where there are $N$ agents and each of them has a private quality drawn independently from a $\mathtt{Bernoulli}(1/N)$ distribution. The quality of an agent represents the value that she can bring to the society. The principal can select up to one agent. If the principal has no additional information, his best strategy is to arbitrarily select an agent to obtain an expected quality of $1/N$. On the other hand, in the “first-best” scenario where the principal has access to full information, he can obtain an expected quality of $1-(1-1/N)^{N}\approx 1-1/e$. The ratio between these values is $\Theta(N)$, which is the worst possible, since random selection provides the same guarantee.
@@
Example 1.2 (continued from Example 1.1).
@@
In the previous setting, one option for all agents is to commit to revealing their true qualities. However, these commitments do not form an equilibrium: If agents $\{2,3,\ldots,N\}$ follow this policy, then agent $1$ is better off not revealing anything, since she can then win with a probability of about $1-1/e\gg 1/N$. In a different scenario where agents $\{2,3,\ldots,N\}$ decide not to disclose anything, agent $1$ can do much better by revealing a quality of 00 with a small probability of $\varepsilon$, and otherwise claiming to have a mean of $\frac{1}{(1-\varepsilon)N}$. Note that this strategy is Bayes plausible, meaning that conditioned on sending the second signal, her posterior probability of being $1$ is indeed $\frac{1}{(1-\varepsilon)N}$. Moreover, she is now selected with probability $1-\varepsilon\gg 1/N$.
@@
The idea of strategic information revelation formalizes voluntary information disclosure in competitive settings and provides a framework for reasoning about the aforementioned real-world scenarios. In this work, we consider a model of competitive Bayesian persuasion first studied by Boleslavsky and Cotton (2015), and subsequently by Au and Kawai (2020) and by Du et al. (2024) among others. In our model, there are multiple agents with private information about their qualities. Those qualities are drawn independently from commonly known priors, which are not necessarily identical. A single principal (or receiver) can select up to $k$ agents and aims at maximizing the expected sum of their qualities. Each agent also derives some utility from being selected based on a given monotone function of her quality. Each agent can disclose information about her quality to the principal via some signaling scheme, which is a probabilistic mapping from qualities to signals. Upon receiving the signals from each agent, the principal updates his belief on each agent’s quality, and finally selects the agents with the highest posterior means of their qualities.
@@
This model formalizes our main goal of characterizing the extent to which competitive information disclosure can improve societal outcomes. As in our example, a natural benchmark for the principal is the “first best”—the sum of qualities of selected agents when the principal has full information of agents’ qualities. As a crucial constraint in Bayesian persuasion models, each agent’s posterior distribution conditioned on their signal must be Bayes plausible, i.e., the signaling scheme must be a valid probabilistic mapping from their qualities to signals. Since agents compete with each other, their overall strategies must form a Nash equilibrium, wherein any agent’s signaling scheme is a best response to all other agents’ strategies based on her own utility function for being selected. Finally, adopting the lens of the price of anarchy (PoA), we define the notion of price of concealment for our setting: How much worse can an equilibrium outcome be, compared to the first-best outcome? Here, we define the price of concealment as the ratio in the principal’s objective (expected sum of qualities of the selected agents) between the full-information outcome and the strategic outcome.
Example 1.3 (continued from Example 1.1).
@@
For $N$ agents with i.i.d. qualities drawn from $\mathtt{Bernoulli}(1/N)$, suppose that each agent gets a constant utility of $1$ from being selected. Then, following the results of Au and Kawai (2020) (see also Theorem 3.1), there is a unique competitive Bayesian persuasion policy in which every agent sends a signal $\sigma_{i}$ (corresponding to their posterior mean) which is statistically identical to an independent sample from a distribution with the cumulative distribution function $G(s)=s^{\frac{1}{N-1}}$. Moreover, the price of concealment is about $2-2/e\approx 1.26$ when $N$ is large.
@@ 1.1 Summary of Results
The main contribution of our work (Theorem 5.1) is to show that the price of concealment of strategic information revelation is a constant (at most $22.19$), even when the agents have different utilities for being selected, the prior distributions over the agents’ qualities are heterogeneous, and the principal can select multiple agents. We also show a lower bound of $2$ on the price of concealment (Theorem 3.1) even in the simple case where the agents’ priors are i.i.d. and Bernoulli.
@@
Before proceeding further, two remarks are in order. First, we define the price of concealment using the principal’s welfare. This definition is somewhat different from the well-known price of anarchy (PoA) definition, as we do not consider the agents’ utility (which can be an arbitrary monotone function of their qualities) in the calculation. [^1] Some of our positive results, in particular, Theorem 4.1, easily extend directly to show the same bound on PoA when we consider the utility of all agents and the principal. This is because in the settings we consider, the principal’s utility—the qualities of the selected agents—matches better with social welfare. In hiring scenarios for example, the agents’ utilities might be derived from receiving monetary payments from the principal, and hence cancel out with the principal’s payments when calculating social welfare. In other scenarios, agents’ utilities need not be on the same scale as each other’s or the principal’s, and might be negligible when compared to the principal’s utility. From now on, we will use the term PoA to refer to our notion (where we consider the gap in the principal’s utility), in keeping with the existing literature.
Next, the work of Au and Kawai (2020) and Du et al. (2024) establishes the existence of an equilibrium in the canonical setting where agents derive constant utility from being selected. [^2] This setting extends the special case presented in Section 4 to non-identical priors. We therefore focus on the quality of the equilibrium as opposed to its existence, a question that has hitherto been unstudied.
@@
Our framework is inspired by techniques from stochastic optimization problems, particularly prophet inequalities (Krengel and Sucheston, 1978) and the core–tail decomposition technique in auction design (Babaioff et al., 2020). We develop related techniques to decompose an agent’s distribution and explicitly construct a signaling scheme that they can use to deviate and improve their utility if the PoA of the equilibrium is too high. This construction is intricate and represents our main technical contribution. As a warm-up to our main result and to present the main ideas of core–tail decomposition and signal construction cleanly, in Section 4, we consider the special case where all agents have identical priors and derive constant utility from being selected. We show a PoA of $4$ for this case and highlight the challenges in generalizing this result.
@@ 1.2 Related Work
Our work is closely related to the field of information design, which studies how information is selectively disclosed to influence decision-making (see surveys by Bergemann and Morris (2019); Dughmi (2017)). A foundational model in this area is Bayesian persuasion (Kamenica and Gentzkow, 2011), where a well-informed sender (in our case, the agent) strategically designs signals to influence the actions of a receiver (in our case, the principal) who seeks to maximize their own utility. This framework has been applied in diverse domains, including price discrimination (Alijani et al., 2022; Bergemann et al., 2015; Banerjee et al., 2024; Cummings et al., 2020; Roesler and Szentes, 2017), security games (Xu et al., 2015), and regret minimization (Babichenko et al., 2021). As mentioned before, the problem has significant non-convexity since the sender does not control the receiver’s optimization routine. Nevertheless, Dughmi and Xu (2016) show computationally efficient algorithms for finding signaling schemes that optimize the sender’s utility. These algorithms do not extend to our setting where there are multiple agents with independent priors. As such, our setting is already challenging even without the strategic aspect, and Banerjee et al. (2025) present logarithmic approximation algorithms for this setting when the receiver chooses one agent.
@@
Unlike classical Bayesian persuasion, where a central mediator controls information flow, our setting features competitive information revelation, where multiple self-interested agents selectively disclose signals about themselves to a decision-maker (receiver or principal). As noted before, an agent’s strategic behavior is in the selection of the signaling scheme, which we assume is Bayes plausible. Special cases of this Bayesian persuasion game have been widely studied in various settings. Boleslavsky and Cotton (2015) analyze how schools strategically set grading standards to influence an evaluator’s perceptions, while Boleslavsky and Cotton (2018) consider project selection under limited capacity, where agents compete by producing persuasive evidence. Similarly, Jain and Whitmeyer (2019) examine competition among senders targeting a rationally inattentive receiver, while Au and Kawai (2021) extend these ideas by incorporating correlated information among senders. These models consider only the competitive behavior between two agents.
@@
For multiple agents, Hwang et al. (2019) study equilibria with competitive information revelation in advertising and pricing, while Hoffmann et al. (2020) investigate the implications of selective disclosure for marketing and privacy regulation. Gradwohl et al. (2022) consider the scenario when the receiver can only obtain signals from a single agent. More recently, Ding et al. (2023) propose a model for competitive information design in a Pandora’s Box framework, and Sapiro-Gheiler (2024) analyze persuasion when receiver preferences are ambiguous. Finally, a different line of work studies the persuasion game when agents share a common state, and reveal information about this state to the receiver (Gentzkow and Kamenica, 2016, 2017; Ravindran and Cui, 2020; Hossain et al., 2024).
@@
In contrast with these models, we study the Bayesian persuasion game with multiple senders and a value-maximizer receiver. The existence of an equilibrium in this model has been shown with symmetric (Au and Kawai, 2020) and asymmetric agents (Du et al., 2024). In their models, every agent gains a constant utility when selected; on the other hand, we study a more general setting where the utilities of each agent can be heterogeneous. Further, our work focuses on the welfare loss arising from strategic signaling via the price of anarchy. As far as we are aware, we are the first to consider the notion of price of anarchy for Bayesian persuasion games.
@@ 2.1 Agent Values and Signaling Schemes
For each agent $i$, her strategy is a signaling scheme $\mathcal{Z}_{i}=\{\Sigma_{i},\pi_{i}\}$, where $\Sigma_{i}$ is some measurable space of signals, and $\pi_{i}\colon[0,1]\to\Delta(\Sigma_{i})$ is a mapping from a realized value $v$ to a probability measure over $\Sigma_{i}$. [^4] Formally, we assume that the function $\pi_{i}$ is a probability kernel from $[0,1]$ to the signal space $\Sigma_{i}$, that is, for every measurable set $S\subseteq\Sigma_{i}$, the mapping $v\mapsto\pi_{i}(S\mid v)$ is measurable. Agent $i$ commits to signaling scheme $\mathcal{Z}_{i}$ before realizing her value; subsequently, after learning $v_{i}$, she samples a signal $\sigma\in\Sigma_{i}$ to send to the principal according to the distribution $\pi_{i}(\cdot\mid v_{i})$.

### v2

@@ 1 Introduction
Every day, people make critical decisions based on limited information, but often someone else controls what information they see. For example, when job seekers apply to a company, they highlight their best qualities (such as a prestigious degree or a successful project) and downplay less flattering information. The employer then must decide who to hire based on the résumés and references provided. Similarly, in college admissions, students choose which recommendation letters to send or whether to submit optional test scores, and hence curate the pictures that admissions officers see. Finally, as a very recent application (Collina et al., 2025), consider an entity that is querying strategic AI agents and using their responses to optimize its action. In all examples, the party presenting the information (for instance, the AI agent or applicant) has an incentive to reveal the information selectively to improve their chances of being selected. Such strategic revelation of information is not lying; it is about being selective in revealing the truth.
From the perspective of someone who aims to optimize societal outcomes, agents revealing information should only be helpful, since the societal optimizer always has the option of ignoring the information revealed. However, quantitatively, to what extent can information revelation improve societal outcomes? We first look at a simple example.
@@
Consider a scenario where there are $N$ agents and each of them has a private quality drawn independently from a $\mathtt{Bernoulli}(1/N)$ distribution. [^1] This means each agent’s quality is $1$ with probability $1/N$, and is 0 with probability 11/N1-1/N. The quality of an agent represents the value that she can bring to the society. The principal can select up to one agent. If the principal has no additional information, his best strategy is to arbitrarily select an agent to obtain an expected quality of $1/N$. On the other hand, in the “first-best” scenario where the principal has access to full information, he can obtain an expected quality of $1-(1-1/N)^{N}\approx 1-1/e$. The ratio between these values is $\Theta(N)$, showing that accessing agents’ value information helps improve the societal outcome significantly.
@@
Example 1.2 (continued from Example˜1.1).
@@
In the previous setting, one option for all agents is to commit to revealing their true qualities. However, these commitments do not form an equilibrium: If agents $\{2,3,\ldots,N\}$ follow this policy, then agent $1$ is better off not revealing anything, since she can then win with a probability of $(1-1/N)^{N-1}\approx 1/e\gg 1/N$. In a different scenario where agents $\{2,3,\ldots,N\}$ decide not to disclose anything, agent $1$ can do much better by sending a signal $\sigma$ with a small probability of $\frac{\varepsilon}{1-1/N}$ conditioned on her value being 0, and otherwise send signal $\sigma^{\prime}$. In this case, the mean of $\sigma^{\prime}$ is $\frac{\Pr[\textrm{her value is $1$}]}{1-\Pr[\sigma\textrm{ is sent}]}=\frac{1/N}{1-\frac{\varepsilon}{1-1/N}\cdot(1-1/N)}=\frac{1}{(1-\varepsilon)N}$. Since $\frac{1}{(1-\varepsilon)N}>1/N$, she will be selected as long as $\sigma^{\prime}$ is sent, with a probability of $1-\varepsilon\gg 1/N$.
@@
The idea of strategic information revelation formalizes voluntary information disclosure in competitive settings and provides a framework for reasoning about the aforementioned real-world scenarios. In this work, we consider a model of competitive Bayesian persuasion first studied by Boleslavsky and Cotton (2015), and subsequently by Au and Kawai (2020) and by Du et al. (2024) among others. In our model, there are multiple agents with private information about their qualities. Those qualities are drawn independently from commonly known priors, which are not necessarily identical. A single principal (or receiver) can select up to $k$ agents and aims to maximize the expected sum of their qualities. Each agent also derives some utility from being selected based on a given monotone function of her quality. Each agent can disclose information about her quality to the principal via some signaling scheme, which is a probabilistic mapping from qualities to signals. Upon receiving the signals from each agent, the principal updates his belief on each agent’s quality, and finally selects the agents with the highest posterior means of quality.
In our model, we assume agents receive utility on selection that is an arbitrary monotone function of their quality. Such monotone utility functions are naturally motivated. As an example, the agents may only care about the event that they are selected, and derive a fixed, short-term utility from selection. This would correspond to the agents’ utility function being a constant independent of their quality. On the other hand, the long-term reward an agent may receive (for instance, pay or grades) subsequent to being selected may depend on their quality. If the agent’s utility depends on such long-term rewards, it can be modeled as a monotone function of quality. Indeed, different agents can prioritize short-term versus long-term rewards differently, leading to heterogeneity in their utility functions.
@@
This signaling model formalizes our main goal of characterizing the extent to which competitive information disclosure can improve societal outcomes. As in our example, a natural benchmark for the principal is the “first best”—the sum of qualities of selected agents when the principal has full information of agents’ qualities. As a crucial constraint in Bayesian persuasion models, each agent’s posterior distribution conditioned on their signal must be Bayes plausible, i.e., the signaling scheme must be a valid probabilistic mapping from their qualities to signals. Since agents compete with each other, their overall strategies must form a Nash equilibrium, wherein any agent’s signaling scheme is a best response to all other agents’ strategies based on her own utility function for being selected. Finally, adopting the lens of the price of anarchy (PoA), we define the notion of price of concealment for our setting: How much worse can an equilibrium outcome be, compared to the first-best outcome? Here, we define the price of concealment as the ratio in the principal’s objective (expected sum of qualities of the selected agents) between the full-information outcome and the strategic outcome.
Example 1.3 (continued from Example˜1.1).
@@
For $N$ agents with i.i.d. qualities drawn from $\mathtt{Bernoulli}(1/N)$, suppose that each agent gets a constant utility of $1$ from being selected. Then, following the results of Au and Kawai (2020) (see also Theorem˜3.1), there is a unique competitive Bayesian persuasion policy in which every agent sends a signal $\sigma_{i}$ (corresponding to their posterior mean) which is statistically identical to an independent sample from a distribution with the cumulative distribution function $G(s)=s^{\frac{1}{N-1}}$. Moreover, the price of concealment is about $2-2/e\approx 1.26$ when $N$ is large.
@@ 1.1 Summary of Results
The main contribution of our work (Theorem˜5.1) is to show that the price of concealment of strategic information revelation is a constant (at most $22.19$), even when the agents have different utilities for being selected, the prior distributions over the agents’ qualities are heterogeneous, and the principal can select multiple agents. We also show a lower bound of $2$ on the price of concealment (Theorem˜3.1) even in the simple case where the agents’ priors are i.i.d. and Bernoulli.
@@
Before proceeding further, three remarks are in order. First, we define the price of concealment using the principal’s welfare. This definition is somewhat different from the well-known price of anarchy (PoA) definition, as we do not consider the agents’ utility (which can be an arbitrary monotone function of their qualities) in the calculation. [^2] Some of our positive results, in particular, Theorem 4.1, easily extend directly to show the same bound on PoA when we consider the utility of all agents and the principal. This is because in the settings we consider, the principal’s utility—the qualities of the selected agents—matches better with social welfare. In hiring scenarios for example, the agents’ utilities might be derived from receiving monetary payments from the principal, and hence cancel out with the principal’s payments when calculating social welfare. In other scenarios, agents’ utilities need not be on the same scale as each other’s or the principal’s, and might be negligible when compared to the principal’s utility. From now on, we will use the term PoA to refer to our notion (where we consider the gap in the principal’s utility), in keeping with the existing literature.
Second, we highlight two fundamental modeling choices that our main result relies on. First, we assume that the agents’ qualities are non-negative, which is natural in settings like hiring or admissions: candidates can be stronger or weaker, but are generally not “harmful” to the principal. As we show in Section˜3, the PoA can become unbounded if qualities are allowed to be negative. The non-negativity also implies that the principal always selects as many agents as allowed. Second, we assume the agents’ utility functions are monotone in their qualities. This assumption ensures that agents’ incentives are not in conflict with the principal’s objective, and deviating toward a signal with more probability mass on higher values will generally increase the agent’s utility. Such property is critical in our utility gain analysis for the deviation signal in Section˜5.3.3.
Finally, the work of Au and Kawai (2020) and Du et al. (2024) establishes the existence of an equilibrium in the canonical setting where agents derive constant utility from being selected, and have identical and non-identical priors respectively. [^3] This setting extends the special case presented in Section 4 to non-identical priors. Though the existence of a Nash Equilibrium is an open question for the general scenario with heterogeneous utility functions that we consider [^4] The equilibrium need not exist for priors with point masses (Du et al., 2024), but it is open whether it exists when the priors are continuous., our constant PoA bound directly applies to the constant utility setting. To circumvent the possible lack of an equilibrium in the general game, we present two natural variants of the game in the appendix, where a Nash equilibrium is guaranteed to exist and the constant PoA bound for the original game applies with minor changes in analysis.
  • In Appendix˜A, we present a discretized variant of the game where the number of signaling schemes, viewed as pure strategies, is finite. Here, under general utilities, we establish existence of Nash equilibria, and show that our PoA results gracefully degrade in the discretization parameter.
  • In Appendix˜B, we introduce a variant where the principal observes the signals with a small random noise, so that their selection is noisy. Under smoothness assumptions on the noise, we show a Nash equilibrium exists, and for sufficiently small noise, our PoA bounds carry over.
Since games are typically implemented in practice via finite menus with discretized reports, and the receiver usually cannot observe signals perfectly, these results indicate that our PoA analysis can be extended to these more realistic, yet general settings, where an equilibrium is guaranteed to exist.
@@
Our framework is inspired by techniques from stochastic optimization problems, particularly prophet inequalities (Krengel and Sucheston, 1978) and the core–tail decomposition technique in auction design (Babaioff et al., 2020). We develop related techniques to decompose an agent’s distribution and explicitly construct a signaling scheme that they can use to deviate and improve their utility if the PoA of the equilibrium is too high. This construction is intricate and represents our main technical contribution. As a warm-up to our main result and to present the main ideas of core–tail decomposition and signal construction cleanly, in Section˜4, we consider the special case where all agents have identical priors and derive constant utility from being selected. We show a PoA of $4$ for this case and highlight the challenges in generalizing this result.
@@ 1.2 Related Work
Our work is closely related to the field of information design, which studies how information is selectively disclosed to influence decision-making (see surveys by Bergemann and Morris (2019); Dughmi (2017)). A foundational model in this area is Bayesian persuasion (Kamenica and Gentzkow, 2011), where a well-informed sender (in our case, the agent) strategically designs signals to influence the actions of a receiver (in our case, the principal) who seeks to maximize their own utility. This framework has been applied in diverse domains, including price discrimination (Alijani et al., 2022; Bergemann et al., 2015; Banerjee et al., 2024; Cummings et al., 2020; Roesler and Szentes, 2017), security games (Xu et al., 2015), AI alignment (Collina et al., 2025), and regret minimization (Babichenko et al., 2021). As mentioned before, the problem has significant non-convexity since the sender does not control the receiver’s optimization routine. Nevertheless, Dughmi and Xu (2016) show computationally efficient algorithms for finding signaling schemes that optimize the sender’s utility. These algorithms do not extend to our setting where there are multiple agents with independent priors. As such, our setting is already challenging even without the strategic aspect, and Banerjee et al. (2025) present logarithmic approximation algorithms for this setting when the receiver chooses one agent.
@@
Unlike classical Bayesian persuasion, where a central mediator controls information flow, our setting features competitive information revelation, where multiple self-interested agents selectively disclose signals about themselves to a decision-maker (receiver or principal). As noted before, an agent’s strategic behavior is in the selection of the signaling scheme, which we assume is Bayes plausible. Special cases of this Bayesian persuasion game have been widely studied in various settings. Boleslavsky and Cotton (2015) analyze how schools strategically set grading standards to influence an evaluator’s perceptions, while Boleslavsky and Cotton (2018) consider project selection under limited capacity, where agents compete by producing persuasive evidence. Similarly, Jain and Whitmeyer (2019) examine competition among senders targeting a rationally inattentive receiver, while Au and Kawai (2021) extend these ideas by incorporating correlated information among senders. Koessler et al. (2022) study a dynamic variant of information disclosure where two competing senders repeatedly release public signals to a single receiver over an unbounded time horizon. They show that disclosure equilibrium is one-shot if there is no constraint on the senders’ disclosure policies. Banerjee et al. (2021) study the price of anarchy in a signaling game, where the agents are restricted to use only threshold tests as signaling strategies. All these models consider only the competitive behavior between two agents.
@@
For multiple agents, Hwang et al. (2019) study equilibria with competitive information revelation in advertising and pricing, while Hoffmann et al. (2020) investigate the implications of selective disclosure for marketing and privacy regulation. Gradwohl et al. (2022) consider the scenario when the receiver can only obtain signals from a single agent. More recently, Ding et al. (2023) propose a model for competitive information design in a Pandora’s Box framework, and Sapiro-Gheiler (2024) analyze persuasion when receiver preferences are ambiguous. Tang et al. (2024) consider a Bayesian persuasion version of prophet inequalities, where the receiver sets thresholds based on known priors of the agents – this converts the game to a two-agent setting, as now each sender is competing only against the receiver, and not other senders. Relatedly, there are other sequential persuasion models inspired by classical stopping problems. Hahn et al. (2020) study Bayesian persuasion through the lens of prophet inequalities, and Hahn et al. (2022) study the secretary recommendation problem in a similar framework. These works focus on sequential arrivals and stopping rules, whereas our work utilizes the prophet-inequality style analysis on Price of Anarchy in a multi-agent competitive environment. Finally, a different line of work studies the persuasion game when agents share a common state, and reveal information about this state to the receiver (Gentzkow and Kamenica, 2016, 2017; Ravindran and Cui, 2020; Hossain et al., 2024).
@@
In contrast with these models, we study the Bayesian persuasion game with multiple senders and a value-maximizer receiver. The existence of an equilibrium in this model has been shown with symmetric (Au and Kawai, 2020) and asymmetric agents (Du et al., 2024). In their models, every agent gains a constant utility when selected; on the other hand, we study a more general setting where the utilities of each agent can be heterogeneous. Further, our work focuses on the welfare loss arising from strategic signaling via the price of anarchy. As far as we are aware, we are the first to consider the notion of price of anarchy for general Bayesian persuasion games.
@@ 2.1 Agent Values and Signaling Schemes
For each agent $i$, her strategy is a signaling scheme $\mathcal{Z}_{i}=(\Sigma_{i},\pi_{i})$, where $\Sigma_{i}$ is some measurable space of signals, and $\pi_{i}\colon[0,1]\to\Delta(\Sigma_{i})$ is a mapping from a realized value $v$ to a probability measure over $\Sigma_{i}$. [^7] Formally, we assume that the function $\pi_{i}$ is a probability kernel from $[0,1]$ to the signal space $\Sigma_{i}$, that is, for every measurable set $S\subseteq\Sigma_{i}$, the mapping $v\mapsto\pi_{i}(S\mid v)$ is measurable. Agent $i$ commits to signaling scheme $\mathcal{Z}_{i}$ before realizing her value; subsequently, after learning $v_{i}$, she samples a signal $\sigma\in\Sigma_{i}$ to send to the principal according to the distribution $\pi_{i}(\cdot\mid v_{i})$.